OpenStudy (anonymous):

suppose you invest some money that grows to the amount A(t)=1000*2^(t/10) in t years a. how much money did you invest? b. how long does it take to double your money?

OpenStudy (anonymous):

\[A(t)=1000(2)^{t/10}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a) The money that first invested is that at t=0, that's: \[A(0)=1000(2)^{0/10}=1000\]

OpenStudy (anonymous):

b)To double this amount, that's to make it 2000: \[1000(2)^{t/10}=2000 \implies 2^{t/10}=2 \implies {t \over 10}=1 \implies t=10\] So, the money will be doubles after 10 years.. Long time I know :P

OpenStudy (anonymous):

why do you use a 0 in the t spot

OpenStudy (anonymous):

Because part a) is asking about the money that you invested at the "beginning", that's at time t=0. So, I plug t=0. Does that make sense?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You're welcome!!